### v1

# AlGaN photonic device platform

Abstract
In the rapidly evolving area of integrated photonics, there is a growing need for materials to support the fast development of advanced and more complex optical on-chip systems. Present photonic material platforms have significantly progressed over the past years; however, at the same time, they face critical challenges. We introduce a novel material for integrated photonics and investigate Aluminum Gallium Nitride (AlGaN) on Aluminum Nitride (AlN) templates as a platform for developing reconfigurable and on-chip nonlinear optical devices. AlGaN combines compatibility with standard fabrication technologies and high electro-optic modulation capabilities with reduced loss in a broad spectrum, making it a viable material for advanced photonic applications. We engineer and grow light-guiding photonic layers and fabricate waveguides, directional couplers, and other on-chip devices. Our AlGaN ring resonators demonstrate tunability and high Q-factors due to the low-loss waveguiding. The comprehensive platform paves the way for nonlinear photon-pair generation applications, on-chip nonlinear quantum frequency conversion, and fast electro-optic modulation for switching and routing.
I Introduction
Advanced classical and quantum photonic applications, such as photonic neuromorphic computing [1], quantum sensing [2] and quantum networking [3, 4], rely on photonic integrated circuits that enable the compact, efficient and high-rate implementation of a variety of optical functionalities. In addition to functionalities adapted from stand-alone optical devices, integrated circuits provide access to mode-multiplexing or routing [5, 6] and efficient fiber coupling. A comprehensive photonic device platform [7], therefore, integrates various components, such as low-loss waveguides, efficient directional couplers, spectral filters, and tunable Mach-Zehnder interferometers (Fig. 1). Another versatile component is the tunable ring resonator, essential for spectral filtering, providing strong light-matter interaction, and enabling nonlinear conversion and photon-pair generation [8, 9, 10]. Finally, due to the large variety of classical and optical components available, the heterogeneous integration of components not fabricated directly into the platform material, such as detectors [11, 12] or quantum light sources [13], is required. A photonic platform material should exhibit low optical losses for efficient light transmission, fast electro-optic modulation capabilities for effectively manipulating optical modes, and significant optical nonlinearities to enable advanced functionalities.
Currently, various material platforms are being explored, each presenting distinct strengths and weaknesses in meeting the comprehensive requirements of a photonic platform. For the ultraviolet and visible spectra, there are several more or less established candidates as an alternative to silicon, which is not transparent in this range. Lithium niobate on insulator is one of the leading platforms due to its outstanding electro-optic coefficient, high optical nonlinearity, and recent progress in technological development [14, 15, 16, 17]. However, it faces several challenges, including the need for non-standard etching processes to fabricate optical components. These processes are incompatible with the conventional manufacturing standards, potentially impeding its scalability [18]. Moreover, despite the advances in epitaxial techniques, the industry still predominantly relies on smart-cut lithium niobate derived from bulk crystals, which typically suffer from optical damage. The limited thickness range (300 nm) of smart-cut lithium niobate restricts some applications, such as nanoscale integrated optics [19]. Silicon nitride (SiN), another established material platform, offers low losses [20] and is compatible with silicon manufacturing standards. SiN-based active optical elements, however, such as tunable MZIs, typically rely on thermo-optic modulation due to their limited electro-optic functionality [21]. Thermo-optic control is comparably slow and generally limited to kHz modulation rates [22]. Another nitride-based material, aluminum nitride (AlN) on sapphire (AlNoS), has been explored due to its relatively high electro-optic constant, its substantial nonlinear coefficient, and its broad transparency, and moderate electro-optic coefficient [23, 24, 25, 26, 27]. AlN on sapphire, however, has yet other limitations. It is grown by sputtering or metallorganic chemical vapor deposition (MOCVD), typically on a sapphire substrate. Consequently, the AlN layer suffers from a high defect density at the sapphire interface [28, 29, 30], leading to relatively high optical scattering and absorption losses. Furthermore, commercially available AlN suffers from significant thickness variation, making it challenging for the scalable fabrication of photonic components.
In this study, we introduce a platform for photonic integrated circuits based on aluminum gallium nitride (AlGaN), where the AlGaN layer acts as a guiding photonic layer that can be patterned into various components. This approach addresses challenges in AlN on sapphire platforms, such as crystal defects at the sapphire-AlN interface. We reduce optical losses and improve device performance by the AlGaN layer, shifting the optical mode away from this interface. The ability to control the ratio of aluminum to gallium in the ternary alloy AlGaN allows for tailored properties of the photonic layer, such as bandgap and refractive index engineering, for specific applications.
Moreover, AlGaN’s wide band gap enables efficient light transmission across a broad spectrum, from UV-C ( 250 nm) to long-wave infrared. Doping AlGaN for electrical conductivity combines optical and electronic functionalities, enabling, for example, current-induced refractive index change [31, 32].
Our AlGaN platform benefits from the advanced state-of-the-art of GaN, which has, over the past decade, evolved into a foundational material for electronic and electro-optic devices, including power electronics and light generation [33, 34]. GaN LEDs and lasers are widely used for producing blue and green light [35, 36]. The epitaxial growth of GaN/AlGaN/InGaN heterostructures on sapphire templates, a well-established practice, opens up possibilities for integrating light sources and electronics on the photonic platform.
The inherent properties of AlGaN, such as its relatively high electro-optic coefficient, make it an ideal candidate for rapid modulation and the development of reconfigurable optical devices. Its significant second-order optical nonlinearity paves the way for on-chip nonlinear photonic devices, including parametric oscillators and sum or difference frequency generators. The generation of correlated photon pairs in the ultraviolet (UV) and visible spectrum has been proposed using an AlGaN/AlN integrated photonics platform via spontaneous four-wave mixing (SFWM) in an AlGaN microring resonator [37]. AlGaN has also been explored as a potential platform for stimulated Brillouin scattering devices such as a racetrack Brillouin laser [38].
Only very few studies have utilized AlGaN as a photonic material. Li et al. [39] fabricated large, multimode waveguides with an AlN/GaN multiple quantum well core and AlGaN claddings. Although this is a notable achievement, the large size and sub-optimal sidewall roughness do not enable application for single-mode photonics. In addition, the wavelength range is limited to around 350 nm and longer. Bruch et al. [40] fabricated ring resonators on AlGaN nanomembranes transferred onto SiO/Si substrates. While their work demonstrated the versatility of AlGaN, using nanomembranes introduces difficulties with handling and scalability.
Here, we develop heterogeneously grown AlGaN into a photonic platform. To fabricate photonic devices, we grow AlGaN, with , directly on high-temperature annealed AlN-on-sapphire templates with a reduced dislocation density[41]. Into this heteromaterial, we fabricate the fundamental constituents of photonic integrated circuits with electron beam lithography: waveguides, Mach-Zehnder interferometers, directional couplers, ring resonators, and fiber couplers.
II Wafer Growth and Characterization
The core ingredient of our novel integrated photonics platform is the heterostructured AlGaN material (Fig. 2(b)). The AlN/AlGaN layer stack was epitaxially grown on a c-plane-oriented sapphire substrate with an offcut of 0.25° towards an m-plane. First, a 350 nm thick AlN layer is deposited using epitaxial magnetron sputtering. The sputtered material’s threading dislocation density (TDD) was decreased through high-temperature annealing (HTA), adhering to the process described by Miyake et al. [42]. During HTA, a temperature of 1700°C is maintained for a relatively short duration of 1h to prevent the formation of aluminum oxynitride on the AlN surface [28]. Subsequently, the rough sputtered and annealed surface is smoothened by metal-organic vapor phase epitaxy (MOVPE) growth of 50 nm AlN in a step flow growth regime. The AlN layer, with a total thickness of 400 nm, exhibited a TDD of , as estimated from high-resolution X-ray diffraction (HRXRD) measurements of the symmetric 0002 and skew-symmetric 10-12 -rocking-curves [43]. Atomic force microscopy revealed a surface root mean square roughness of 0.09 nm over a area.
Next, a 250 nm thick AlGaN waveguide layer is grown by MOVPE in a 6 x 2-inch close-coupled showerhead reactor on the AlN template, utilizing trimethylaluminum, triethylgallium, and ammonia as source materials, with hydrogen as the carrier gas, and a growth temperature of 1015°C. In-situ reflectometry measurements verify the attainment of the targeted layer thickness of 250 nm. After AlGaN growth, we measure the surface RMS roughness to be 0.5 nm. This smooth surface supports integrating photonic devices as introduced here and enables the hetero-integration of superconducting nanowire single-photon detectors (SNSPDs) [11] and quantum devices. The strain relaxation towards the underlying AlN buffer and the Al-mole fraction x of the layers is determined by HRXRD using reciprocal space maps (RSM) of the 0004 and 1124 reflections. Each RSM shows two sharp peaks (Fig. 2(a)), corresponding to the AlN template and AlGaN layers. By analyzing these data, an Al-mole fraction x of 0.69 and a relaxation of the in-plane compressive strain of only 3 can be determined for the AlGaN layer. This shows that the AlGaN grew almost pseudomorphically on the AlN template, and the generation of misfit dislocations for strain relief can thus be excluded.
III Optical Material Properties
We used the prism coupling technique as described in [44, 45] to quantify the optical losses in the as-grown thin films. It utilizes a rutile prism to couple a tunable white laser (SuperK Fianium) into the films. The losses are quantified by monitoring the decay of the scattered light with a digital camera. We characterize a 400 nm AlNoS sample and an AlGaN/AlN/sapphire wafer. The AlNoS supports a single lateral TE mode at 630 nm, whereas the AlGaN wafer exhibits two distinct modes, TE and TE, at the same wavelength. By fine-tuning the coupling angle, we isolate and analyze the losses of these modes, as reported in Fig. 2(d).
The TE mode, sustained by the AlGaN wafer, exhibits optical losses that decrease with increasing wavelength, reaching a minimum value of 0.37 cm above 600 nm. In contrast, the TE mode of the AlGaN and the TE mode of the AlNoS exhibited higher losses, with minimum values of approximately 1.36 and 1.0 cm, respectively, to rise again for wavelengths over 625 nm. It is worth noting that both of these modes are localized at the interface between two different materials, and as such, they are sensitive to the interface’s inherent roughness. Additional sources of loss could be due to defects and color centers in AlN and are presumably concentrated at the AlN-sapphire interface as shown by TEM measurements by [28, 29, 30].
AlN has defects comprising a mix of impurities, vacancies, and lattice defects. Factors like threading dislocations influence the formation of these defects, although the exact conditions for their formation are yet not fully understood [46, 47, 48],
IV Device Fabrication
In the fabrication process of waveguides and ring resonators, we initially deposit a 200 nm thick SiO layer using Plasma Enhanced Chemical Vapor Deposition (PECVD) to serve as a hard mask. We then coat a 10 nm thick Ti layer via e-beam evaporation to act as a charge dissipation layer, followed by a 400 nm ZEP 520A e-beam lithography resist. The process steps are illustrated in Fig. 3(a). After patterning the resist, we transfer the pattern to the hard mask using a two-step reactive ion etching process: we utilize SF plasma to etch Ti and CF plasma to etch SiO. In the final etching step, we process the AlGaN layer using BCl:Cl:He at a ratio of 10:50:10 sccm, under 600 W ICP power and 100 W RF power, at a pressure of 1 Pa. The AlGaN to SiO etch rate ratio is approximately 3/1. We implement a laser-cut notch at the chip’s edge to facilitate precise cleaving. Fig. 3(b) presents an SEM image of a cleaved facet, revealing the AlN and AlGaN layers. We measure the waveguides to have a wall angle of 75. Fig. 3(c) shows the simulated power profiles of the fundamental TE mode of the waveguides at wavelengths of 632 nm and 785 nm. At 632 nm, 95, and at 785 nm, 90 of the optical power resides above the AlN-sapphire interface, hence well removed from this lossy boundary.
V Tunable Microring Resonators
We fabricate microring resonators, incorporating radii of 50 m and top widths of 580 nm, and coupled them to a racetrack geometry waveguide, cleaved at both ends, as depicted in Fig. 4(a). The sample is mounted on a 3D piezoelectric nanopositioner (Physik Instrumente, NanoCube) and characterized with a custom-made inverted microscope, where the same objective (Zeiss, LD EC Epiplan-Neofluar, 20X, NA = 0,22) is used to launch the laser light into one waveguide facet and collect it from the other facet. The alignment of the incoupling light is monitored with an EMCCD camera (Andor, iXon Ultra 897) and optimized by maximizing the intensity of the outcoupled signal by adjusting both sample position and laser incoupling angle. It should be noted that the diffraction-limited spot of the laser focused on the waveguide facet is significantly larger than the size of the mode supported by the waveguide. Therefore, we cannot determine the effective incoupled power. Propagation losses are then inferred by measuring the intrinsic quality factor of the microring resonators. To reach critical coupling, which allows for the estimation of resonator losses, we fabricated several microring resonators coupled to racetrack waveguides with varied coupling gaps, ranging from 100-325 nm, as displayed in Fig. 4(c). We measure the transmission of a tunable 785 nm laser (Toptica, DFB PRO centered at 784.6 nm, MHF over 2 nm) and use a linear film polarizer (Thorlabs, LPVIS) to excite and collect TE or TM modes selectively. Fig. 4(b) illustrates the typical transmission as a function of the laser’s frequency, with a baseline primarily attributed to the laser power modulation, which is typical for DFB. This modulation can be described with a quadratic baseline from which the resonant peaks can be clearly discriminated and fitted with a Lorentzian function, as indicated in red in Fig. 4(b). We extrapolated the resonance widths at the critical coupling, i.e., zero transmittance. We summarize key parameters such as the critical coupling FWHM, free spectral ranges (FSRs), and intrinsic Q-factors for TE and TM modes in Table 1.
We mount the chips on a custom-made temperature-controlled stage to test the temperature tunability of the TE and TM modes resonance frequencies in the range of 20-30 C. Fig. 4(c) depicts the shift of a resonant peak of the TE mode (coupling gap=100nm) for different temperatures. In the investigated temperature range, the central frequency of the resonance changes linearly with temperature. The corresponding temperature coefficients are 5.7 and 6.2 GHz/K for TE and TM modes. These coefficients correspond to a thermo-optic coefficient for AlGaN, which is comparable to the previously reported values for AlN (2.94 10 K) and GaN (7.01 10 K) [49] at 785 nm and 300K. We also determine the anisotropic refractive indices of the AlN and AlGaN layers by analyzing the FSRs of rings with different widths and etch depths, as explained in the supplementary material. The thermo-optic coefficient and effective index results for TM and TE modes are summarized in Table 1.
| Parameter | TE | TM |
| FWHM at Critical Coupling (GHz) | 2.5 | 7.0 |
| FSR (GHz) | 438 | 426 |
| Propagation loss (cm) | 0.57 | 1.643 |
| Intrinsic Q factor | ||
| at 785 nm (GHz/K) | 5.7 | 6.2 |
| 2.9 | 3.26 | |
| at 632 nm | 2.14 | 2.20 |
| at 632 nm | 2.07 | 2.11 |
VI Directional Couplers
Directional couplers are essential components in photonic integrated circuits. They enable the transfer of electromagnetic energy between two or more waveguides that are in proximity. They can be combined to form optical switches or Mach-Zehnder interferometers, which are the fundamental constituent of the majority of on-chip quantum technology applications [50, 51]. In our study, we develop AlGaN directional couplers linked to racetrack waveguides, as depicted in Fig. 5(a). We vary coupling gaps and lengths, as illustrated in Fig. 5(b). By injecting a 632 nm laser into one of the input ports, we measure the output intensity ratio: . Our observations revealed a direct linear correlation between the coupling length and , with linear fitting of the data providing the coupling constants. For coupling gaps of 230 nm and 330 nm, the coupling constants were determined to be 0.15 cm and 0.066 cm , respectively. The observed uncertainties are primarily attributed to reflections at the input and output facets, which can be minimized by applying an anti-reflection coating. The well-functioning directional couplers show that our AlGaN platform is well-suitable for building the required components of complex integrated photonic circuits. Further improvement is to be expected by optimizing the etching processes towards vertical sidewalls.
VII Discussion
Our investigation of AlGaN as a material platform for photonic integrated circuits allowed us to demonstrate the required core functional devices. By epitaxially growing AlGaN on HTA-AlN templates on sapphire substrates, we alleviate the problems related to crystal defects and dislocations at the sapphire-AlN interface. Our measured losses align with those documented in the literature concerning aluminum nitride on sapphire waveguides [52, 24, 53], even though we did not employ a cladding material, which is expected to diminish scattering loss even more. Additionally, we observed a polarization rotation effect within the rings, which might effectively increase the measured loss; however, this could be mitigated by optimizing waveguide geometry. Moreover, by measuring the thermo-optic coefficient and anisotropic refractive indices of the waveguides, we have gained insight into parameters relevant to applications such as nonlinear photon-pair generation and on-chip nonlinear conversion. The birefringent refractive index opens avenues for exploration, especially in its utilization for phase matching of TE and TM modes. Furthermore, the surface roughness of the epitaxial material is sufficiently low to integrate superconducting single-photon detectors. In summary, our work represents a significant step forward in using AlGaN for integrated photonics, offering a scalable and high-quality alternative to established materials.
Acknowledgements.
We thank Karina Ickert, Nico Sluka, Joost Wartena, Adrian Runge, Mohammad Mohammadi, and Dominik Sudau for their assistance in thin-film depositions; Kevin Kunkel, Natalia Sabelfeld, Kai Gehrke, and Andreas Renkewitz for performing plasma etching processes; Alexander Külberg and Felix Eiche for laser cutting and marking; Uwe Spengler and Christine Münnich for cleaving, Olaf Krüger, Ralph-Stephan Unger and Ina Ostermay for fruitful discussions.References
- Shastri et al. [2021] B. J. Shastri, A. N. Tait, T. Ferreira de Lima, W. H. P. Pernice, H. Bhaskaran, C. D. Wright, and P. R. Prucnal, Photonics for artificial intelligence and neuromorphic computing, Nature Photonics 15, 102–114 (2021).
- Xavier et al. [2021] J. Xavier, D. Yu, C. Jones, E. Zossimova, and F. Vollmer, Quantum nanophotonic and nanoplasmonic sensing: towards quantum optical bioscience laboratories on chip, Nanophotonics 10, 1387 (2021).
- Knaut et al. [2023] C. M. Knaut, A. Suleymanzade, Y.-C. Wei, D. R. Assumpcao, P.-J. Stas, Y. Q. Huan, B. Machielse, E. N. Knall, M. Sutula, G. Baranes, N. Sinclair, C. De-Eknamkul, D. S. Levonian, M. K. Bhaskar, H. Park, M. Lončar, and M. D. Lukin, Entanglement of Nanophotonic Quantum Memory Nodes in a Telecommunication Network, (2023).
- Zhang et al. [2019] G. Zhang, J. Y. Haw, H. Cai, F. Xu, S. M. Assad, J. F. Fitzsimons, X. Zhou, Y. Zhang, S. Yu, J. Wu, W. Ser, L. C. Kwek, and A. Q. Liu, An integrated silicon photonic chip platform for continuous-variable quantum key distribution, Nature Photonics 13, 839–842 (2019).
- Liu et al. [2019] Y. Liu, K. Xu, S. Wang, W. Shen, H. Xie, Y. Wang, S. Xiao, Y. Yao, J. Du, Z. He, and Q. Song, Arbitrarily routed mode-division multiplexed photonic circuits for dense integration, Nature Communications 10, 10.1038/s41467-019-11196-8 (2019).
- Zeng et al. [2021] Z. Zeng, L. Lu, P. He, D. Liu, and M. Zhang, Demonstration of dual-mode photonic integrated circuit based on inverse-designed photonic components, IEEE Photonics Technology Letters 33, 1289 (2021).
- Siew et al. [2021] S. Y. Siew, B. Li, F. Gao, H. Y. Zheng, W. Zhang, P. Guo, S. W. Xie, A. Song, B. Dong, L. W. Luo, C. Li, X. Luo, and G.-Q. Lo, Review of silicon photonics technology and platform development, Journal of Lightwave Technology 39, 4374–4389 (2021).
- Logan et al. [2018] A. D. Logan, M. Gould, E. R. Schmidgall, K. Hestroffer, Z. Lin, W. Jin, A. Majumdar, F. Hatami, A. W. Rodriguez, and K.-M. C. Fu, 400W second harmonic conversion efficiency in 14 m-diameter gallium phosphide-on-oxide resonators, Optics Express 26, 33687 (2018).
- Guo et al. [2016] X. Guo, C.-l. Zou, C. Schuck, H. Jung, R. Cheng, and H. X. Tang, Parametric down-conversion photon-pair source on a nanophotonic chip, Light: Science Applications 6, e16249–e16249 (2016).
- Wang et al. [2023] X. Wang, X. Jiao, B. Wang, Y. Liu, X.-P. Xie, M.-Y. Zheng, Q. Zhang, and J.-W. Pan, Quantum frequency conversion and single-photon detection with lithium niobate nanophotonic chips, npj Quantum Information 9, 10.1038/s41534-023-00704-w (2023).
- Zhu et al. [2016] D. Zhu, H. Choi, T.-J. Lu, Q. Zhao, A. Dane, F. Najafi, D. R. Englund, and K. K. Berggren, Superconducting nanowire single-photon detector on aluminum nitride, in 2016 Conference on Lasers and Electro-Optics (CLEO) (2016) pp. 1–2.
- Esmaeil Zadeh et al. [2021] I. Esmaeil Zadeh, J. Chang, J. W. N. Los, S. Gyger, A. W. Elshaari, S. Steinhauer, S. N. Dorenbos, and V. Zwiller, Superconducting nanowire single-photon detectors: A perspective on evolution, state-of-the-art, future developments, and applications, Applied Physics Letters 118, 10.1063/5.0045990 (2021).
- Mouradian et al. [2015] S. L. Mouradian, T. Schröder, C. B. Poitras, L. Li, J. Goldstein, E. H. Chen, M. Walsh, J. Cardenas, M. L. Markham, D. J. Twitchen, M. Lipson, and D. Englund, Scalable Integration of Long-Lived Quantum Memories into a Photonic Circuit, Phys. Rev. X 5, 031009 (2015).
- Poberaj et al. [2012] G. Poberaj, H. Hu, W. Sohler, and P. Günter, Lithium niobate on insulator (LNOI) for micro-photonic devices, Laser Photonics Reviews 6, 488 (2012).
- Qi and Li [2020] Y. Qi and Y. Li, Integrated lithium niobate photonics, Nanophotonics 9, 1287–1320 (2020).
- Saravi et al. [2021] S. Saravi, T. Pertsch, and F. Setzpfandt, Lithium niobate on insulator: An emerging platform for integrated quantum photonics, Advanced Optical Materials 9, 10.1002/adom.202100789 (2021).
- Zhu et al. [2021] D. Zhu, L. Shao, M. Yu, R. Cheng, B. Desiatov, C. J. Xin, Y. Hu, J. Holzgrafe, S. Ghosh, A. Shams-Ansari, E. Puma, N. Sinclair, C. Reimer, M. Zhang, and M. Lončar, Integrated photonics on thin-film lithium niobate, Adv. Opt. Photonics 13, 242 (2021).
- Han et al. [2023] X. Han, M. Yuan, H. Xiao, G. Ren, T. G. Nguyen, A. Boes, Y. Su, A. Mitchell, and Y. Tian, Integrated photonics on the dielectrically loaded lithium niobate on insulator platform, Journal of the Optical Society of America B 40, D26 (2023).
- Zivasatienraj et al. [2021] B. Zivasatienraj, M. B. Tellekamp, and W. A. Doolittle, Epitaxy of LiNbO: Historical challenges and recent success, Crystals 11, 397 (2021).
- Buzaverov et al. [2023] K. A. Buzaverov, A. S. Baburin, E. V. Sergeev, S. S. Avdeev, E. S. Lotkov, M. Andronik, V. E. Stukalova, D. A. Baklykov, I. V. Dyakonov, N. N. Skryabin, M. Y. Saygin, S. P. Kulik, I. A. Ryzhikov, and I. A. Rodionov, Low-loss silicon nitride photonic ICs for near-infrared wavelength bandwidth, Optics Express 31, 16227 (2023).
- Hermans et al. [2019] A. Hermans, M. V. Daele, J. Dendooven, S. Clemmen, C. Detavernier, and R. Baets, Integrated silicon nitride electro-optic modulators with atomic layer deposited overlays, Opt. Lett. 44, 1112 (2019).
- Nejadriahi et al. [2020] H. Nejadriahi, A. Friedman, R. Sharma, S. Pappert, Y. Fainman, and P. Yu, Thermo-optic properties of silicon-rich silicon nitride for on-chip applications, Opt. Express 28, 24951 (2020).
- Li et al. [2021] N. Li, C. P. Ho, S. Zhu, Y. H. Fu, Y. Zhu, and L. Y. T. Lee, Aluminium nitride integrated photonics: a review, Nanophotonics 10, 2347–2387 (2021).
- Sun et al. [2019] Y. Sun, W. Shin, D. A. Laleyan, P. Wang, A. Pandey, X. Liu, Y. Wu, M. Soltani, and Z. Mi, Ultrahigh Q microring resonators using a single-crystal aluminum-nitride-on-sapphire platform, Optics Letters 44, 5679 (2019).
- Wu et al. [2020] X. Wu, J. Feng, X. Liu, , and H. Zeng, Effects of rapid thermal annealing on aluminum nitride waveguides, Optical Materials Express 10, 10.1364/OME.410129 (2020).
- Liu et al. [2018] X. Liu, A. W. Bruch, Z. Gong, J. Lu, J. B. Surya, L. Zhang, J. Wang, J. Yan, and H. X. Tang, Ultra-high-Q UV microring resonators based on a single-crystalline AlN platform, 5, 1279 (2018).
- Liu et al. [2017a] X. Liu, C. Sun, B. Xiong, L. Wang, J. Wang, Y. Han, Z. Hao, H. Li, Y. Luo, J. Yan, T. Wei, Y. Zhang, and J. Wang, Aluminum nitride-on-sapphire platform for integrated high-Q microresonators, Opt. Express 25, 587 (2017a).
- Hagedorn et al. [2019] S. Hagedorn, S. Walde, A. Mogilatenko, M. Weyers, L. Cancellara, M. Albrecht, and D. Jaeger, Stabilization of sputtered AlN/sapphire templates during high temperature annealing, Journal of Crystal Growth 512, 142 (2019).
- Tokumoto et al. [2009] Y. Tokumoto, N. Shibata, T. Mizoguhci, T. Yamamoto, and Y. Ikuhara, Atomic structure of threading dislocations in AlN thin films, Physica B: Condensed Matter 404, 4886 (2009).
- Dogheche et al. [2000] E. Dogheche, B. Belgacem, D. Remiens, P. Ruterana, and F. Omnes, Prism coupling as a non-destructive tool for optical characterization of (Al, Ga) nitride compounds, MRS Internet Journal of Nitride Semiconductor Research 5, 747 (2000).
- Schwarz et al. [2004] U. T. Schwarz, E. Sturm, W. Wegscheider, V. Kümmler, A. Lell, and V. Härle, Excitonic signature in gain and carrier induced change of refractive index spectra of (In, Al)GaN quantum well lasers, Applied Physics Letters 85, 1475 (2004).
- Bulutay et al. [2010] C. Bulutay, C. M. Turgut, and N. A. Zakhleniuk, Carrier-induced refractive index change and optical absorption in wurtzite InN and GaN: Full-band approach, Physical Review B 81, 10.1103/physrevb.81.155206 (2010).
- Udabe et al. [2023] A. Udabe, I. Baraia-Etxaburu, and D. G. Diez, Gallium nitride power devices: A state of the art review, IEEE Access 11, 48628 (2023).
- A. Tabbakh et al. [2023] T. A. Tabbakh, D. Anandan, M. J. Sheldon, P. Tyagi, and A. Alfaifi, Recent advancements in GaN LED technology, in Light-Emitting Diodes - New Perspectives (IntechOpen, 2023).
- Liang et al. [2021] F. Liang, D. Zhao, Z. Liu, P. Chen, J. Yang, L. Duan, Y. Shi, and H. Wang, GaN-based blue laser diode with 6.0 W of output power under continuous-wave operation at room temperature, Journal of Semiconductors 42, 112801 (2021).
- Hu et al. [2020] L. Hu, X. Ren, J. Liu, A. Tian, L. Jiang, S. Huang, W. Zhou, L. Zhang, and H. Yang, High-power hybrid GaN-based green laser diodes with ITO cladding layer, Photonics Research 8, 279 (2020).
- [37] F. D. Leonardis, R. A. Soref, M. Soltani, and V. M. N. Passaro, Broadband biphoton generation and statistics of quantum light in the UV-visible range in an AlGaN microring resonator 10.1038/s41598-017-11617-y.
- Leonardis et al. [2018] F. D. Leonardis, R. A. Soref, M. Soltani, and V. M. Passaro, Stimulated Brillouin scattering in an AlGaN photonics platform operating in the visible spectral range, Scientific Reports 8, 10.1038/S41598-018-33183-7 (2018).
- Li et al. [2004] N. Li, I. Waki, C. Kumtornkittikul, J. H. Liang, M. Sugiyama, Y. Shimogaki, and Y. Nakano, Fabrication of AlGaN-based waveguides by inductively coupled plasma etching, Japanese Journal of Applied Physics, Part 2: Letters 43, 10.1143/JJAP.43.L1340 (2004).
- Bruch et al. [2017] A. W. Bruch, K. Xiong, H. Jung, X. Guo, C. Zhang, J. Han, and H. X. Tang, Electrochemically sliced low loss AlGaN optical microresonators, Applied Physics Letters 110, 10.1063/1.4973521 (2017).
- Hagedorn et al. [2020] S. Hagedorn, S. Walde, A. Knauer, N. Susilo, D. Pacak, L. Cancellara, C. Netzel, A. Mogilatenko, C. Hartmann, T. Wernicke, M. Kneissl, and M. Weyers, Status and prospects of AlN templates on sapphire for ultraviolet light‐emitting diodes, physica status solidi (a) 217, 10.1002/pssa.201901022 (2020).
- Miyake et al. [2016] H. Miyake, C.-H. Lin, K. Tokoro, and K. Hiramatsu, Preparation of high-quality AlN on sapphire by high-temperature face-to-face annealing, Journal of Crystal Growth 456, 155 (2016).
- Pantha et al. [2007] B. N. Pantha, R. Dahal, M. L. Nakarmi, N. Nepal, J. Li, J. Y. Lin, H. X. Jiang, Q. S. Paduano, and D. Weyburne, Correlation between optoelectronic and structural properties and epilayer thickness of AlN, Applied Physics Letters 90, 10.1063/1.2747662 (2007).
- Weber et al. [1973] H. P. Weber, F. A. Dunn, and W. N. Leibolt, Loss measurements in thin-film optical waveguides, Appl. Opt. 12, 755 (1973).
- Dogheche et al. [1999] E. Dogheche, D. Rémiens, A. Boudrioua, and J. C. Loulergue, Growth and optical characterization of aluminum nitride thin films deposited on silicon by radio-frequency sputtering, Applied Physics Letters 74, 1209 (1999).
- Cannon et al. [2023] J. K. Cannon, S. G. Bishop, J. P. Hadden, H. B. Yağcı, and A. J. Bennett, Polarization study of single color centers in aluminum nitride, Applied Physics Letters 122, 10.1063/5.0145542 (2023).
- Varley et al. [2016] J. B. Varley, A. Janotti, and C. G. V. de Walle, Defects in AlN as candidates for solid-state qubits, Physical Review B 93, 10.1103/physrevb.93.161201 (2016).
- Xue et al. [2022] Y. Xue, J. Yuan, Q. Li, F. Chen, X. Yuan, Z. Ju, S. Zhang, B. Wu, Y. Hou, M. Li, J. Zhang, and E. Wu, Deciphering the photophysical properties of near-infrared quantum emitters in AlGaN films by transition dynamics, Nanoscale 14, 18115 (2022).
- Watanabe et al. [2008] N. Watanabe, T. Kimoto, and J. Suda, The temperature dependence of the refractive indices of GaN and AlN from room temperature up to 515 °C, Journal of Applied Physics 104, 10.1063/1.3021148 (2008).
- Zhou et al. [2022] H. Zhou, J. Dong, J. Cheng, W. Dong, C. Huang, Y. Shen, Q. Zhang, M. Gu, C. Qian, H. Chen, Z. Ruan, and X. Zhang, Photonic matrix multiplication lights up photonic accelerator and beyond, Light: Science Applications 11, 10.1038/s41377-022-00717-8 (2022).
- Clements et al. [2016] W. R. Clements, P. C. Humphreys, B. J. Metcalf, W. S. Kolthammer, and I. A. Walmsley, Optimal design for universal multiport interferometers, Optica 3, 1460 (2016).
- Liu et al. [2017b] Y. Liu, Q. X. Li, L. Y. Wan, B. Kucukgok, E. Ghafari, I. T. Ferguson, X. Zhang, S. Wang, Z. C. Feng, and N. Lu, Composition and temperature dependent optical properties of AlGaN alloy by spectroscopic ellipsometry, Applied Surface Science 421, 389 (2017b).
- Pernice et al. [2012] W. H. Pernice, C. Xiong, C. Schuck, and H. X. Tang, Second harmonic generation in phase matched aluminum nitride waveguides and micro-ring resonators, Applied Physics Letters 100, 10.1063/1.4722941 (2012).